# On the interaction of dilatancy and friction in the behavior of fluid–saturated sheared granular materials: a coupled Computational Fluid Dynamics–Discrete Element Method study

### v1

@@ VI.1 Pore pressure evolution during breaching
The evolution of excess pore pressure during breaching reveals distinct responses in dense and loose subaqueous granular systems, governed by the initial packing density and fluid–particle interactions. Fig. 9 presents the spatial and temporal variations across monitored points A–L (A–F for the dense model shown in Fig. 9(a), and G–L for the loose model shown in Fig. 9(d)), which were strategically selected to capture localized pressure dynamics.
@@
In the dense model, wall removal initiates an abrupt drop in pore pressure, followed by a gradual rise as fluid redistributes and particles rearrange (Fig. 9(b)). This transient suction phase results from the system’s resistance to compaction. Over time, pore pressure stabilizes, reflecting the balance between fluid expulsion and interparticle contact forces. At point D, the temporal evolution of excess pore pressure (Fig. 9(c)), plotted in Lagrangian coordinates across multiple time intervals (0.01 s, 0.1 s, 0.4 s, and 0.8 s), shows an initial negative pressure that progressively increases toward a steady state. This behavior aligns well with analytical predictions, with minor deviations during the transient phase attributed to localized heterogeneities and nonlinear fluid–particle coupling.
@@
In contrast, the loose model exhibits a sharp rise in pore pressure following wall removal, driven by void collapse and rapid fluid compression (Fig. 9(e)). The high initial pressure reflects the system’s susceptibility to compaction, followed by redistribution and a gradual decline toward equilibrium. At point J, the temporal pressure profile (Fig. 9(f)) begins elevated and stabilizes over time. While the steady-state matches analytical solutions well, significant deviations are observed during the transient stage, underscoring the complex dynamics of loosely packed systems.
@@
These comparisons confirm that the analytical solution, despite its simplifying assumptions, effectively captures the steady-state behavior of excess pore pressure for both dense and loose subaqueous breaching scenarios. The good agreement observed at later times across different monitored points supports the applicability of the poromechanical framework in predicting long-term pore pressure evolution. While transient mismatches highlight the complexity of fluid–particle interactions during early deformation, the steady-state comparison underscores the robustness of the analytical model in capturing the essential physics governing pore pressure regulation during breaching.
@@ VII Conclusion
This study investigated the collapse and runout behavior of granular materials in subaerial (dry) and subaqueous (fluid-saturated) environments, with a particular focus on the roles of initial packing density and fluid–particle interactions. Three-dimensional CFD-DEM simulations revealed that dense packings form narrower shear zones and exhibit reduced runout distances due to higher resistance to deformation. In contrast, loose packings undergo broader deformation and longer runouts as a result of easier compaction.
@@
In subaqueous conditions, these trends are amplified: dense materials fail gradually through dilation and the generation of negative pore pressures, whereas loose materials experience rapid failure driven by compaction and transient positive pore pressures. The evolution of porosity plays a central role in controlling deformation behavior, fluid migration, and mechanical stability. Dense systems dilate during shear, drawing in fluid and increasing resistance, while loose systems compact, expelling fluid and decreasing stability. These behaviors are governed by the strain rate and local stress conditions. The granular rheology of subaerial and subaqueous models under dense and loose conditions further provides critical insights into the interplay between friction, strain rate, and dimensionless parameters such as the inertial number ($I_{n}$) and viscous number ($I_{v}$). Both environments exhibit a general rate-strengthening behavior, though driven by distinct mechanisms: in subaerial systems, friction increases modestly with $I_{n}$ due to enhanced contact networks and force-chain persistence, while in subaqueous systems, frictional evolution is influenced by the competing effects of fluid drag, pore-pressure buildup, and lubrication. Dense materials tend to maintain nearly constant porosity at low $I_{n}$ and $I_{v}$, reflecting a dynamic balance between compaction and dilation, whereas loose materials exhibit more pronounced porosity changes and frictional variability at higher strain rates. These rheological trends highlight the sensitivity of granular strength to both fluid presence and initial structure.
@@
Simulation results were consistent with analytical predictions of excess pore pressure, demonstrating the model’s ability to capture the complex coupling between fluid flow and granular deformation at steady state, while transient mismatches highlighted the complexity of fluid–particle interactions during early deformation. Overall, the study demonstrates that granular failure dynamics are highly sensitive to initial structure and hydromechanical interactions. These findings contribute to a deeper understanding of natural phenomena such as landslides, sediment transport, and seabed instabilities. To further enhance predictive capability, future work should incorporate time-dependent effects, spatial heterogeneity in material properties, and comparison against field-scale observations. In addition, to better capture the stick-slip behavior associated with frictional evolution, we plan to derive rate-and-state parameters from small-scale direct shear experiments. These parameters will then be integrated into a frictional rheology model based on the rate-and-state framework, enabling more accurate representation of transient shear responses in fluid-saturated granular systems.
@@ VIII Supplementary Material
Supplementary material is available to support the findings presented in this study. It provides additional figures and explanatory content detailing the poromechanical behavior of the modeled granular systems. Included materials cover the temporal evolution of permeability, derivation of isotropic unloading compressibility and dilation strength parameters, and boundary conditions applied for analytical model comparison. These supplemental results offer further insight into the influence of confining pressure, pore structure, and unloading on fluid migration and material response during breaching.
@@ Acknowledgments
The research has been supported by start-up funds from the Department of Civil and Environmental Engineering and the Cullen College of Engineering at the Univeristy of Houston, as well as the American Chemical Society (ACS) Petroleum Research Fund under ACS-PRF Grant 66470-DNI8.
@@
The MFIX simulation scripts and open-source code used in this study are available on Zenodo under the title: “MFIX Simulation Scripts and Open Source Code for research paper entitled ’On the interaction of dilatancy and friction in the behavior of fluid-saturated sheared granular materials: a coupled Computational Fluid Dynamics–Discrete Element Method study”. The dataset can be accessed at: https://doi.org/10.5281/zenodo.16415581.

### v2

@@ VI.1 Pore pressure evolution during breaching
The evolution of excess pore pressure during breaching reveals distinct responses in dense and loose subaqueous granular systems, governed by the initial packing density and fluid–particle interactions. Fig. 15 presents the spatial and temporal variations across monitored points (A, B, D, E, and F for the dense configuration shown in Fig. 15(a), and G, H, J, K, and L for the loose configuration shown in Fig. 15(d)), which were strategically selected to capture localized pressure dynamics.
@@
In the dense configuration, wall removal initiates an abrupt drop in pore pressure, followed by a gradual rise as fluid redistributes and particles rearrange (Fig. 15(b)). This transient suction phase results from the system’s resistance to compaction. Over time, pore pressure stabilizes, reflecting the balance between fluid expulsion and interparticle contact forces. At point D, the temporal evolution of excess pore pressure (Fig. 15(c)), plotted in Lagrangian coordinates across multiple time intervals (0.01 s, 0.1 s, 0.4 s, and 0.8 s), shows an initial negative pressure that progressively increases toward a steady state. This behavior aligns well with analytical predictions, with minor deviations during the transient phase attributed to localized heterogeneities and nonlinear fluid–particle coupling.
@@
In contrast, the loose configuration exhibits a sharp rise in pore pressure following wall removal, driven by void collapse and rapid fluid compression (Fig. 15(e)). The high initial pressure reflects the system’s susceptibility to compaction, followed by redistribution and a gradual decline toward equilibrium. At point J, the temporal pressure profile (Fig. 15(f)) begins elevated and stabilizes over time. While the steady-state matches analytical solutions well, significant deviations are observed during the transient stage, underscoring the complex dynamics of loosely packed systems.
@@
These comparisons confirm that the analytical solution, despite its simplifying assumptions, effectively captures the steady-state behavior of excess pore pressure for both dense and loose subaqueous breaching scenarios. The good agreement observed at later times across different monitored points supports the applicability of the poromechanical framework in predicting long-term pore pressure evolution. The transient mismatches highlight the complexity of fluid–particle interactions during early deformation, and the possible role of transient frictional behavior of the granular materials in response to perturbations kohler2023rate; grelle2010shear; lemos2004shear.
The observation that transient pore pressure does not align with the analytical soil-mechanics model during the initial flow phases can inform further development of submerged granular-flow models, such as the compressible $\mu(J)$, $\phi(J)$ formulation proposed by rauter2021compressible. Rauter’s model provides a compressible two-phase continuum description of submerged granular motion by coupling the Navier–Stokes equations with a $\mu(J)$, $\phi(J)$ rheology. The constitutive relations $\mu(J)$ and $\phi(J)$ incorporate dilatancy, pore-pressure coupling, and critical-state behavior, offering a continuum representation of dense, submerged granular flow. Incorporating quantitative relationships derived from our rheological fitting and the transient pore-pressure responses observed in our CFD–DEM simulations could extend Rauter’s model to capture the poromechanical feedbacks that control flow initiation, and transient particle–fluid interactions.
@@ VII Conclusion
This study investigated the collapse and runout of granular materials in subaerial (dry) and subaqueous (fluid-saturated) settings, focusing on the roles of initial packing density and fluid–particle coupling. Three-dimensional CFD–DEM simulations showed that dense packings formed narrower shear zones and shorter runouts due to greater resistance to deformation, whereas loose packings deformed more broadly and ran out farther because compaction is easier.
@@
In subaqueous conditions, these trends were amplified: dense materials failed gradually via dilation and the generation of negative pore pressures, while loose materials failed rapidly due to compaction and transient positive pore pressures. Porosity evolution controlled deformation, fluid migration, and mechanical stability. Dense systems dilated during shear, drawing in fluid and increasing resistance; loose systems compacted, expelling fluid and reducing stability. These behaviors were governed by the strain rate and local stress state. Rheologically, the subaerial and subaqueous models under dense and loose conditions revealed how friction, strain rate, and the inertial ($I_{n}$) and viscous ($I_{v}$) numbers interact. Both environments exhibited overall rate strengthening, but for different reasons: in subaerial flows, friction increased modestly with $I_{n}$ due to enhanced contact networks and force-chain persistence, whereas in subaqueous flows, frictional evolution reflected competing effects of fluid drag, pore-pressure buildup, and lubrication. Dense materials tended to maintain nearly constant porosity at low $I_{n}$ and $I_{v}$, indicating a dynamic balance between compaction and dilation, while loose materials showed larger porosity changes and greater frictional variability at higher strain rates. These trends highlight the sensitivity of granular strength to both fluid presence and initial packing fraction.
@@
Simulation results agree with analytical predictions of excess pore pressure as flow approaches steady state, demonstrating that the model captures the coupling between fluid flow and granular deformation. Early-time discrepancies, however, underscore the complexity of fluid–particle interactions during transients. Overall, granular failure dynamics are highly sensitive to initial structure and hydromechanical coupling. These findings inform efforts to mitigate natural hazards such as landslides, sediment transport, and seabed instabilities. To enhance predictive capability, future work should incorporate time-dependent effects, spatial heterogeneity in material properties, and systematic comparison with field-scale observations. Moreover, because transient frictional behavior in granular materials may contribute to the early-time mismatch between analytical predictions and the CFD–DEM evolution of pore pressure, future models should explicitly include transient (rate-and-state) friction and be evaluated against analytical solutions that incorporate this mechanism.
@@ VIII Supplementary Material
See the supplementary material for additional figures and explanatory content detailing the poromechanical behavior of the modeled granular systems. Included materials cover the temporal evolution of permeability, derivation of isotropic unloading compressibility and dilation strength parameters, and boundary conditions applied for analytical model comparison. These supplemental results offer further insight into the influence of confining pressure, pore structure, and unloading on fluid migration and material response during breaching.
@@ Acknowledgments
The research has been supported by start-up funds from the Department of Civil and Environmental Engineering and the Cullen College of Engineering at the University of Houston, as well as the American Chemical Society (ACS) Petroleum Research Fund under ACS-PRF Grant 66470-DNI8. We also thank Prof. Ruth Cardinaels, Associate Editor of Physics of Fluids, and the anonymous reviewers for their constructive feedback, which improved the clarity and rigor of this manuscript.
@@
The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.16415581, Ref. ferdowsi_2025_16415581.